3. For the following A.P.s, write the first term and the common difference.
(i) 3, 1, – 1, – 3 …
(ii) -5, – 1, 3, 7 …
(iii) 1/3, 5/3, 9/3, 13/3 ….
(iv) 0.6, 1.7, 2.8, 3.9 …

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solutions (i) Given series, 3, 1, – 1, – 3 … First term, a = 3 Common difference, d = Second term – First term ⇒ 1 – 3 = -2 ⇒ d = -2 (ii) Given series, – 5, – 1, 3, 7 … First term, a = -5 Common difference, d = Second term – First term ⇒ ( – 1)-( – 5) = – 1+5 = 4 (iii) Given series, 1/3, 5/3, 9/3, 13/3 …. First term, a = 1/3 Common difference, d = Second term – First term ⇒ 5/3 – 1/3 = 4/3 (iv) Given series, 0.6, 1.7, 2.8, 3.9 … First term, a = 0.6 Common difference, d = Second term – First term ⇒ 1.7 – 0.6 ⇒ 1.1